OpenStudy (anonymous):

Algebra 2 help please!!??

OpenStudy (anonymous):

Both \(x\) and \(y\) increase together, thus this is direct variation. Inverse variation would mean, only one variable increase while the other may stay constant or decrease. Can you tell me the formula for direct variation?

OpenStudy (anonymous):

i don't know it..

OpenStudy (anonymous):

Well the formula is: \[y=kx\] Remember that: \[k \ne 0 \text{ and k represents the constant of variation.}\] Lets use the table. Notice the \(x\) and \(y\) values. Plug in the values: \[9 = k(3) \implies 9 \div 3 = k \implies 3 = k\]

OpenStudy (anonymous):

3=k?

OpenStudy (anonymous):

Yes, \(k = 3\) stands true for all the values :-). Now do you know how to write a function that models/represents this?

OpenStudy (anonymous):

no..

OpenStudy (anonymous):

can you show me how?:)

OpenStudy (anonymous):

Okay! I presume you have heard of \(y = mx + c\). (C maybe B depending where you live.) \(y = mx + c\) is a function. Since we know \(k\) is the constant variation, we are allowed to write it as: \[f(x) = 3x~\text{or}~y=3x\]

OpenStudy (anonymous):

You can learn more about functions from: http://www.purplemath.com/modules/fcnnot.htm http://www.khanacademy.org/math/algebra/algebra-functions/classic-function-videos/v/introduction-to-functions It is really hard to explain in just text, normally you learn about functions before you do direct variation and inverse variation. Anyway we have answered your question, that was all there was to it, easy right? :-)